The percentage by mole of NO2​(g) in a mixture of NO2​(g) and NO(g) having average molecular mass 34 is :

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
Dear student

Dear Student,

 

Lets assume that the mole fraction of NO2 is ‘x’ and NO is ‘y’.

 

Now lets assume that the container contains only these two gases. So we can say that,

 

x+y=1

 

Now we know that molecular weight of NO2 is 46 and that of NO is 30. Now,the avg molecular mass is given by

 

M1.x +M2.y=34

 

where M1 and M2 are the individual molecular masses

 

Now solving the two equations we find that , x= 0.25 and y= 0.75

 

Percentage by mole of NO2 = x * 100 = 0.25*100 = 25%

 

regards

team askiitians